OpenStudy (anonymous):

Describe the locus of points whose coordinates satisfy both statements: x2 = 49 x2 + y2 = 85

OpenStudy (anonymous):

x^2+y^2=r^2 7^2+y^2=85 y^2=85-49 y=√36 y=±6

OpenStudy (anonymous):

What next?

OpenStudy (anonymous):

is the first equation only x^2=49?? no y component??

OpenStudy (anonymous):

so u get only four points on a circle with radius sqrt(85) the points are (+-7,+-6)

OpenStudy (anonymous):

u can't find locus here only u can find the solution...

OpenStudy (anonymous):

this is what the thing on my worksheet says. :(

OpenStudy (anonymous):

x=+-7 and y= +-6

OpenStudy (anonymous):

how did you find x±7?

OpenStudy (anonymous):

just satisfy the eqn x=+-7 with x^2 = 49

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(7,6), (-7,6), (7,-6),(-7,-6)

OpenStudy (anonymous):

all the points above satisfy your equation....They are vertices of a rectangle ;)

OpenStudy (anonymous):

thank you!!!